# A NICER View of PSR J1231--1411: A Complex Case

### v1

@@ 3.3.1 Mass and Inclination Priors
We use mass $M$ and inclination $i$ priors that are based on the preliminary Shapiro delay measurement results from Cromartie et al. (2024), which were obtained through a Bayesian timing model fit to data from the Robert C. Byrd Green Bank Telescope and from the NRT. To approximate the 2D probability density function, we first fitted a Gaussian function to the marginalized probability distribution of $M$ using a subset of the Markov Chain Monte Carlo samples from the Shapiro delay analysis (those having $M$ between 1 to 3 $M_{\odot}$). We obtained a mean of $\sim 1.002$ $M_{\odot}$ and a standard deviation of $\sim 0.930$ $M_{\odot}$. We inverse sampled the mass from this distribution (truncating it between 1 and 3 $M_{\odot}$).  [^9] We note that the resulting prior is still quite close to uniform. After that, we inverse sampled the $\cos i$ from another Gaussian, whose mean and standard deviations were set to depend on the value of the already drawn mass; $\mu_{\cos i}(M)$, $\sigma_{\cos i}(M)$. The values were found by fitting Gaussians to the $\cos i$ probability distributions for a grid of masses and by fitting their dependence on mass with a polynomial function. This resulted in the following:
@@
The resulting 2D prior for $M$ and $\cos i$ is shown in Figure 3. We see that highest probabilities are at the smallest allowed masses but that the distribution fades very slowly towards higher masses. Inclination is slightly better constrained, peaking above $80$ degrees ($\cos i\lesssim 0.18$). We note that the final prior adopted in X-PSI is additionally modified by the compactness and surface gravity conditions (as e.g., in Riley et al. 2021) to prevent the polar radius from being smaller than the photon sphere ($R_{\textrm{polar}}/r_{\rm g}(M)>3$, where $r_{\rm g}$ is the Schwarzchild radius), and to not extrapolate intensities beyond the precomputed atmosphere grid (e.g., $13.7\leq\log_{10}g\leq 15.0$, for the fully-ionized hydrogen grid, where $g$ is the surface gravity). This makes the prior probability at smaller masses even higher since fewer NS samples fail the compactness condition there, hence more are retained.
Figure 3: Two-dimensional $\cos i-M$ prior distribution based on the radio timing measurements from Cromartie et al. (2024). The color indicates the probability of a $\cos i-M$ cell.
@@ 3.3.2 Radius Prior
For the NS (equatorial) radius, X-PSI analyses have usually assumed a prior that is initially flat between $\sim 4.4$ km (i.e., $3r_{\rm g}$ for a $1$ $M_{\odot}$ star) and $16$ km, but then modified by the compactness and surface gravity conditions mentioned also in Section 3.3.1. In this work, we apply also more informative radius priors. In particular, we consider four cases when choosing a radius prior. In the first scenario, we use the standard X-PSI radius prior. In the second scenario, we set the lower limit of radius to be $8$ km to avoid extremely compact solutions. In the third scenario, we restrict the radius to be between $10$ and $14$ km. In the fourth scenario (called “R21”), we require the radius to be consistent with previous observational and theoretical EOS constraints by using the posterior samples from Raaijmakers et al. (2021) [^10] https://zenodo.org/doi/10.5281/zenodo.4696231 to numerically build up a one-dimensional prior distribution for radius. We selected the samples from the “combined constraints” case obtained with the “PP” parametrization of the EOS corresponding roughly a Gaussian centered around 12 km (see the lower left panel in Figure 5 of Raaijmakers et al. 2021). We converted the samples to a cumulative probability distribution using the numpy.cumsum function and interpolated from that distribution with scipy.interpolate.Akima1DInterpolator when sampling.
@@ 3.3.3 Other Priors
For all the other model parameters we use fairly uninformative priors. For the geometry parameters describing the hot regions we follow the choices of Riley et al. (2021); Vinciguerra et al. (2024), i.e., we have initially uniform priors in the spherical cap angular radii (limited to be below $\pi/2$), in their phase coordinates, and in cosines of their colatitudes. These priors are then modified by the requirements that primary and secondary hot regions cannot overlap (but in case of PDT-U the two caps forming a hot region must overlap), and that the primary hot region is always the one with smaller co-latitude. The hot region temperatures are uniform in $\log_{10}T_{\mathrm{eff}}$ and bound based on the limits of the precomputed atmosphere model grid (e.g., $\log_{10}T_{\mathrm{eff}}\in[5.1,6.8]$ for fully-ionized hydrogen). When including blackbody emission from the rest of the star’s surface, we sample its temperature $\log_{10}{T_{\mathrm{else}}}$ uniformly between $5.0$ and $6.5$ as in Vinciguerra et al. (2024).
Unlike for the previous NSs analyzed with NICER, the distance $D$ to PSR J1231--1411 is known rather poorly. Yao et al. (2017) estimated the $D$ to be $420$ pc using the dispersion measure of PSR J1231--1411 and an electron density model of the Galaxy. However, they also reported a distance range between $350$ and $510$ pc based on independent estimates from optical measurements of the white dwarf companion. In this work, we use slightly more conservative limits and sample $D$ uniformly between $100$ and $700$ pc. We note, though, that preliminary parallax distance measurements indicate a distance of roughly $600\pm 100$ pc (Cromartie et al., 2024).
Since the interstellar hydrogen column density $N_{\mathrm{H}}$ is also not well known, we sample that uniformly between 00 and $10^{21}\mathrm{cm}^{-2}$ as in Riley et al. (2021). In addition, the effective-area scaling factor priors are those described already in Section 3.1.
@@ 3.4 Likelihood and Background
We construct a joint NICER and XMM-Newton likelihood function and marginalize the likelihoods over background variables (one for each detector channel) in the same way as described in Section 2.5 of Riley et al. (2021). We do not directly impose any informative limits on the NICER background, but for XMM-Newton instruments we bound the background marginalization to be done within $\mathscr{B}_{\rm X}\pm 4\sqrt{\mathscr{B}_{\rm X}}$ (but always above 00), where $\mathscr{B}_{\rm X}$ is the measured number of background counts for a given channel (see the dashed lines in Figure 2 after being rescaled based on the BACKSCAL factor, i.e., the difference in the extraction area between the PSR J1231--1411 and background observation). The prior in background is assumed to be uniform between the lower and upper limits. Constraining the background using XMM-Newton offers also an indirect constraint on the NICER background by restricting the number of source photons that can satisfy both NICER and XMM-Newton data simultaneously. In the end, we cross-check the inferred NICER background with that predicted by the SCORPEON background model for NICER (see Section 2.1 and 4.2).
@@ 3.5 Posterior Computation
We compute the posterior samples using PyMultiNest (Buchner et al., 2014) and MultiNest (Feroz & Hobson, 2008; Feroz et al., 2009, 2019), as in the previous X-PSI analyses. In all runs we set both the sampling efficiency (defined as described in Appendix B of Salmi et al. 2024) and the evidence tolerance to $0.1$. We also keep the multimode/mode-separation setting of MultiNest turned on to get more information about the multiple modes found in the posterior volume of PSR J1231--1411. We checked for one ST-U test run that turning the multimode setting off had no significant effects on the results or on the computation time. For ST-U runs we used 4,000 live points, which was deemed sufficient based on preliminary test runs with older data and mass-inclination prior versions. For PDT-U runs we increased the number of live points to 10,000 and to 20,000. However, these were found to provide likely converged results only in the case of the two most restricted radius priors (the third and fourth cases described in Section 3.3.2). Due to the already high computational cost (above 200,000 core-hours for a 20,000 live point run), better convergence by increasing the live points further or by dropping the sampling efficiency was not attempted.
@@
Figure 4: Bolometric pulse profiles (counts summed over energy channels) for the maximum likelihood parameters (red stepped curves) of the ST-U H $T_{\mathrm{else}}$ (upper left panel) and PDT-U H $R_{\mathrm{eq}}\in[10,14]$ km (upper right panel) km models. See Table 1 and Section 4.2 for model explanations. The bolometric NICER data are shown with black stepped curves. The differences (data -- model) are shown in the bottom panels also with black stepped curves. The expected Poisson fluctuation levels (for NICER data) are shown there with blue dashed curves. We note that the model curves are rather similar between all the ST-U versions and between all the PDT-U versions.
@@
Figure 5: Radius, mass, compactness and inclination posterior distributions using the NICER and XMM-Newton data sets for the PDT-U model. The results with three different radius priors are shown, but the run with widest radius prior ($R_{\mathrm{eq}}\in[8,16]$ km) seems not converged as it finds only significantly worse fits to the data than the others. Dash-dotted curves represent the marginal prior probability density functions (PDFs). The vertical dashed orange line shows $R_{\mathrm{eq}}=14$ km to guide the eye. The shaded vertical bands show the $68.3\%$ credible intervals (for the posteriors corresponding to the red curves), and the contours in the off-diagonal panels show the $68.3\%$, $95.4\%$, and $99.7\%$ credible regions. See the captions of Figure 5 of Riley et al. (2021) for additional details about the figure elements.
@@
Figure 6: The NICER count data (top), posterior-expected count numbers (middle, averaged from 200 equally weighted posterior samples), and (Poisson) residuals (bottom) for the PDT-U H $R_{\mathrm{eq}}\in[10,14]$ km model. See Figure 6 of Riley et al. (2021) for additional details about the figure elements. The complete figure set (7 images), which includes the residuals for all of the results presented in Figure 5, is available in the online journal (HTML version).
@@
Note. —   The evidence ($\widehat{\ln\mathcal{Z}}$) and maximum likelihood ($\widehat{\ln\textrm{ML}}$) differences are shown compared to the PDT-U He $R_{\mathrm{eq}}$ R21 model. We also show the bolometric NICER $\chi^{2}$ values ($\chi^{2}_{\textrm{ML,bol}}$) corresponding the maximum likelihood sample. Note that the exact d.o.f. is not known, but it is something smaller than $32$ (see Footnote 11). For ST-U, only runs with the broadest radius prior are considered here; besides the standard case (ST-U H), values are also shown when including emission from rest of the star surface (ST-U H $T_{\mathrm{else}}$) and when having 3 different XMM-Newton effective-area scaling factors (ST-U H $3\alpha$).
Figure 7: Inferred NICER background for the joint NICER and XMM-Newton analysis using the PDT-U H $R_{\mathrm{eq}}\in[10,14]$ km model. The blue solid stepped curve shows the total NICER count-rate spectrum. The orange stepped curves show the backgrounds that maximize the likelihood for $1000$ equally weighted posterior samples. The red curve shows the corresponding background for the maximum likelihood sample. The black dashed (dot-dashed) curve shows the highest (lowest) background when considering the model uncertainty as explained in Section 2.1. The inferred backgrounds of the other presented models are very similar to those shown here.
Figure 8: Posterior-expected pulse profiles for the joint NICER and XMM-Newton analysis using the PDT-U H $R_{\mathrm{eq}}\in[10,14]$ km model. The top and top-center panels show the signal incident on the telescopes and bottom-center and bottom panels show the signal registered by NICER after accounting for the instrument response. The blue dashed curves show the posterior-expected signals generated by the two hot regions separately and the black solid curves in combination. The conditional posterior distribution of the incident photon flux and the NICER count rate are also presented (purple contours in the top-center and bottom panels, respectively). See Figure 11 of Riley et al. (2021) for additional details about the figure elements.
Figure 9: The maximum likelihood geometry configuration for the joint NICER and XMM-Newton analysis using the PDT-U H $R_{\mathrm{eq}}\in[10,14]$ km model. The viewing angle represents the Earth inclination to the spin axis. The maximum likelihood geometry is very similar for the PDT-U H $R_{\mathrm{eq}}$ R21 and PDT-U He $R_{\mathrm{eq}}$ R21 models, except that the primary superseding component is located at the equator for the latter.
@@ 4.1 ST-U
We start with the inference for PSR J1231--1411 using the ST-U model. We applied the model using three different versions of it: (1) Basic ST-U, (2) ST-U where the rest of the star’s surface is also emitting (as mentioned in Section 3.2), (3) ST-U where all 3 XMM-Newton instruments have their effective-area scaling factors as free parameters (as mentioned in Section 3.1). The corresponding posteriors for the spacetime parameters are shown in Appendix A for the standard X-PSI radius prior (posteriors for the other parameters are available in the Zenodo, Salmi et al. 2024). We found that in all three cases the inferred radius is very small, around $5$ km, and the NS is extremely compact, even though the mass is only slightly above $1$ $M_{\odot}$. In fact, all these solutions strongly hit our prior upper bound on surface gravity set by the hydrogen atmosphere table limits ($\log_{10}g=15.0$).
While no very obvious patterns were observed in the residuals between the inferred phase-energy resolved signals and the PSR J1231--1411 data (which would indicate a poor fit to the data), we found that the fit to the phase-averaged (or bolometric) NICER pulse profile is rather bad for all the ST-U models (see the left panel of Figure 4 for an example). The bolometric $\chi^{2}$ values for maximum likelihood solutions are around $80$ for $32$ phase bins [^11] The degree of freedom (d.o.f.) is not obvious, but it should be a few smaller than $32$ depending on how many of the model parameters influence the bolometric pulse profile (Dittmann et al., 2024). We also note that better bolometric $\chi^{2}$ values are found when looking at the $1000$ highest likelihood samples in each ST-U run, since the best bolometric NICER $\chi^{2}$ value does not exactly coincide with the maximum joint (2D) NICER and XMM-Newton likelihood sample. However, these $\chi^{2}$ values are still always above $70$.. In addition, if forcing the radius to be above $8$ km (making a full run with the second radius prior case mentioned in Section 3.3.2), both the maximum likelihood and the corresponding bolometric $\chi^{2}$ values became significantly worse: The former dropped by more than 20 in $\ln$-units and the latter increased to $\sim 160$ (results of this run are available in the Zenodo, Salmi et al. 2024). Most notably, the best-fit model was no longer able to produce the interpulse around phase $0.5$ that is seen in the data (see Figure 1). For the high compactness star this was easier due to multiple images of the hot spots contributing to the interpulse. The highest Bayesian evidence for a ST-U model (the one including rest of the star emission) was also found to be significantly smaller than for any of the PDT-U models (by $8-33$ in $\ln$-units), so we therefore focus on the latter next.

### v2

@@ 3.3.1 Mass and Inclination Priors
We use mass $M$ and inclination $i$ priors that are based on the preliminary Shapiro delay measurement results from Cromartie et al. (2024, in preparation), which were obtained through a Bayesian timing model fit to data from the Robert C. Byrd Green Bank Telescope and from the NRT. To approximate the 2D probability density function, we first fitted a Gaussian function to the marginalized probability distribution of $M$ using a subset of the Markov chain Monte Carlo samples from the Shapiro delay analysis (those having $M$ between 1 and 3 $M_{\odot}$). We obtained a mean of $\sim 1.002$ $M_{\odot}$ and a standard deviation of $\sim 0.930$ $M_{\odot}$. We inverse sampled the mass from this distribution (truncating it between 1 and 3 $M_{\odot}$).  [^23] We note that the resulting prior is still quite close to uniform. After that, we inverse sampled the $\cos i$ from another Gaussian, whose mean and standard deviations were set to depend on the value of the already drawn mass; $\mu_{\cos i}(M)$, $\sigma_{\cos i}(M)$. The values were found by fitting Gaussians to the $\cos i$ probability distributions for a grid of masses and by fitting their dependence on mass with a polynomial function. This resulted in the following:
@@
The resulting 2D prior for $M$ and $\cos i$ is shown in Figure 3. We see that highest probabilities are at the smallest allowed masses but that the distribution fades very slowly toward higher masses. Inclination is slightly better constrained, peaking above $80\arcdeg$ ($\cos i\lesssim 0.18$). We note that the final prior adopted in X-PSI is additionally modified by the compactness and surface gravity conditions (as e.g., in Riley et al. 2021) to prevent the polar radius from being smaller than the photon sphere ($R_{\textrm{polar}}/r_{\rm g}(M)>3$, where $r_{\rm g}$ is the Schwarzchild radius), and to not extrapolate intensities beyond the precomputed atmosphere grid (e.g., $13.7\leq\log_{10}g\leq 15.0$, for the fully ionized hydrogen grid, where $g$ is the surface gravity). This makes the prior probability at smaller masses even higher because fewer NS samples fail the compactness condition there, hence more are retained.
Figure 3: A 2D $\cos i-M$ prior distribution based on the radio timing measurements from Cromartie et al. (2024, in preparation). The color indicates the probability of a $\cos i-M$ cell.
@@ 3.3.2 Radius Prior
For the NS (equatorial) radius, X-PSI analyses have usually assumed a prior that is initially flat between $\sim 4.4$ km (i.e., $3r_{\rm g}$ for a $1$ $M_{\odot}$ star) and $16$ km, but then modified by the compactness and surface gravity conditions, mentioned also in Section 3.3.1. In this work, we apply also more informative radius priors. In particular, we consider four cases when choosing a radius prior. In the first scenario, we use the standard X-PSI radius prior. In the second scenario, we set the lower limit of radius to be $8$ km to avoid extremely compact solutions. In the third scenario, we restrict the radius to be between $10$ and $14$ km. In the fourth scenario (called “R21”), we require the radius to be consistent with previous observational and theoretical EOS constraints by using the posterior samples from Raaijmakers et al. (2021, on Zenodo, ) to numerically build up a 1D prior distribution for radius. We selected the samples from the “combined constraints” case obtained with the “PP” parametrization of the EOS corresponding roughly a Gaussian centered around 12 km (see the lower left-hand panel in Figure 5 of Raaijmakers et al. 2021). We converted the samples to a cumulative probability distribution using the numpy.cumsum function and interpolated from that distribution with scipy.interpolate.Akima1DInterpolator when sampling.
@@ 3.3.3 Other Priors
For all the other model parameters, we use fairly uninformative priors. For the geometry parameters describing the hot regions, we follow the choices of Riley et al. (2021) and Vinciguerra et al. (2024), i.e., we have initially uniform priors in the spherical cap angular radii (limited to be below $\pi/2$), in their phase coordinates and in cosines of their colatitudes. These priors are then modified by the requirements that primary and secondary hot regions cannot overlap (but in case of PDT-U the two caps forming a hot region must overlap), and that the primary hot region is always the one with smaller colatitude. The hot region temperatures are uniform in $\log_{10}T_{\mathrm{eff}}$ and bound based on the limits of the precomputed atmosphere model grid (e.g., $\log_{10}T_{\mathrm{eff}}\in[5.1,6.8]$ for fully ionized hydrogen). When including blackbody emission from the rest of the star’s surface, we sample its temperature $\log_{10}{T_{\mathrm{else}}}$ uniformly between $5.0$ and $6.5$ as in Vinciguerra et al. (2024).
Unlike for the previous NSs analyzed with NICER, the distance $D$ to PSR J1231--1411 is known rather poorly. Yao et al. (2017) estimated the $D$ to be $420$ pc using the DM of PSR J1231--1411 and an electron density model of the Galaxy. However, they also reported a distance range between $350$ and $510$ pc based on independent estimates from optical measurements of the white dwarf companion. In this work, we use slightly more conservative limits and sample $D$ uniformly between $100$ and $700$ pc. We note, though, that preliminary parallax distance measurements indicate a distance of roughly $600\pm 100$ pc (Cromartie et al. 2024, in preparation).
Since the interstellar hydrogen column density $N_{\mathrm{H}}$ is also not well known, we sample that uniformly between 00 and $10^{21}\mathrm{cm}^{-2}$, as in Riley et al. (2021). In addition, the effective-area scaling factor priors are those described already in Section 3.1.
@@ 3.4 Likelihood and Background
We construct a joint NICER and XMM-Newton likelihood function and marginalize the likelihoods over background variables (one for each detector channel) in the same way as described in Section 2.5 of Riley et al. (2021). We do not directly impose any informative limits on the NICER background, but for XMM-Newton instruments we bound the background marginalization to be done within $\mathscr{B}_{\rm X}\pm 4\sqrt{\mathscr{B}_{\rm X}}$ (but always above 00), where $\mathscr{B}_{\rm X}$ is the measured number of background counts for a given channel (see the dashed lines in Figure 2 after being rescaled based on the BACKSCAL factor, i.e., the difference in the extraction area between the PSR J1231--1411 and background observation). The prior in the background is assumed to be uniform between the lower and upper limits. Constraining the background using XMM-Newton also offers an indirect constraint on the NICER background by restricting the number of source photons that can satisfy both NICER and XMM-Newton data simultaneously. In the end, we cross-check the inferred NICER background with that predicted by the SCORPEON background model for NICER (see Sections 2.1 and 4.2).
@@ 3.5 Posterior Computation
We compute the posterior samples using PyMultiNest (Buchner et al., 2014) and MultiNest (Feroz & Hobson, 2008; Feroz et al., 2009, 2019), as in the previous X-PSI analyses. In all runs, we set both the sampling efficiency (defined as described in Appendix B of Salmi et al. 2024) and the evidence tolerance to $0.1$. We also keep the multimode/mode-separation setting of MultiNest turned on to get more information about the multiple modes found in the posterior volume of PSR J1231--1411. We checked for one ST-U test run that turning the multimode setting off had no significant effects on the results or on the computation time. For ST-U runs, we used 4000 live points, which was deemed sufficient based on preliminary test runs with older data and mass-inclination prior versions. For PDT-U runs, we increased the number of live points to 10,000 and to 20,000. However, these were found to provide likely converged results only in the case of the two most restricted radius priors (the third and fourth cases described in Section 3.3.2). Due to the already high computational cost (above 200,000 core hours for a 20,000 live point run), better convergence by increasing the live points further or by dropping the sampling efficiency was not attempted.
@@
Figure 4: Bolometric pulse profiles (counts summed over energy channels) for the maximum-likelihood parameters (red-stepped curves) of the ST-U H $T_{\mathrm{else}}$ (upper left-hand panel) and PDT-U H $R_{\mathrm{eq}}\in[10,14]$ km (upper right-hand panel) km models. See Table 1 and Section 4.2 for model explanations. The bolometric NICER data are shown with black-stepped curves. The differences (data -- model) are shown in the bottom panels also with black-stepped curves. The expected Poisson fluctuation levels (for NICER data) are shown there with blue-dashed curves. We note that the model curves are rather similar between all the ST-U versions and between all the PDT-U versions.
@@
Figure 5: Radius, mass, compactness, and inclination posterior distributions using the NICER and XMM-Newton data sets for the PDT-U model. The results with three different radius priors are shown, but the run with widest radius prior ($R_{\mathrm{eq}}\in[8,16]$ km) seems not to have converged because it finds only significantly worse fits to the data than the others. Dashed-dotted curves represent the marginal prior probability density functions (PDFs). The vertical dashed orange line shows $R_{\mathrm{eq}}=14$ km to guide the eye. The shaded vertical bands show the $68.3\%$ credible intervals (for the posteriors corresponding to the red curves), and the contours in the off-diagonal panels show the $68.3\%$, $95.4\%$, and $99.7\%$ credible regions. See the captions of Figure 5 of Riley et al. (2021) for additional details about the figure elements.
@@
Figure 6: The NICER count data (top panel), posterior-expected count numbers (middle panel, averaged from 200 equally weighted posterior samples), and (Poisson) residuals (bottom panel) for the PDT-U H $R_{\mathrm{eq}}\in[10,14]$ km model. See Figure 6 of Riley et al. (2021) for additional details about the figure elements. The complete figure set (seven images), which includes the residuals for all of the models listed in Table 1, is available in the online journal (HTML version).
@@
Note. —   The evidence ($\widehat{\ln\mathcal{Z}}$) and maximum-likelihood ($\widehat{\ln\textrm{ML}}$) differences are shown compared to the PDT-U He $R_{\mathrm{eq}}$ R21 model. We also show the bolometric NICER $\chi^{2}$ values ($\chi^{2}_{\textrm{ML,bol}}$) corresponding the maximum-likelihood sample. Note that the exact d.o.f. is not known, but it is something smaller than $32$ (see Footnote 24). For ST-U, only runs with the broadest radius prior are considered here; besides the standard case (ST-U H), values are also shown when including emission from rest of the star surface (ST-U H $T_{\mathrm{else}}$) and when having three different XMM-Newton effective-area scaling factors (ST-U H $3\alpha$).
Figure 7: Inferred NICER background for the joint NICER and XMM-Newton analysis using the PDT-U H $R_{\mathrm{eq}}\in[10,14]$ km model. The blue solid stepped curve shows the total NICER count-rate spectrum. The orange stepped curves show the backgrounds that maximize the likelihood for $1000$ equally weighted posterior samples. The red curve shows the corresponding background for the maximum-likelihood sample. The black-dashed (dotted-dashed) curve shows the highest (lowest) background when considering the model uncertainty, as explained in Section 2.1. The inferred backgrounds of the other presented models are very similar to those shown here.
Figure 8: Posterior-expected pulse profiles for the joint NICER and XMM-Newton analysis using the PDT-U H $R_{\mathrm{eq}}\in[10,14]$ km model. The top and top-center panels show the signal incident on the telescopes, and the bottom-center and bottom panels show the signal registered by NICER after accounting for the instrument response. The blue-dashed curves show the posterior-expected signals generated by the two hot regions separately and the black-solid curves in combination. The conditional posterior distribution of the incident photon flux and the NICER count rate are also presented (purple contours in the top-center and bottom panels, respectively). See Figure 11 of Riley et al. (2021) for additional details about the figure elements.
Figure 9: The maximum-likelihood geometry configuration for the joint NICER and XMM-Newton analysis using the PDT-U H $R_{\mathrm{eq}}\in[10,14]$ km model. The viewing angle represents the Earth’s inclination to the spin axis. The maximum-likelihood geometry is very similar for the PDT-U H $R_{\mathrm{eq}}$ R21 and PDT-U He $R_{\mathrm{eq}}$ R21 models, except that the primary superseding component is located at the equator for the latter.
@@ 4.1 ST-U
We start with the inference for PSR J1231--1411 using the ST-U model. We applied the model using three different versions of it: (1) Basic ST-U, (2) ST-U where the rest of the star’s surface is also emitting (as mentioned in Section 3.2), (3) ST-U where all three XMM-Newton instruments have their effective-area scaling factors as free parameters (as mentioned in Section 3.1). The corresponding posteriors for the spacetime parameters are shown in the Appendix A for the standard X-PSI radius prior (posteriors for the other parameters are available in the Zenodo, Salmi et al. 2024 (doi:10.5281/zenodo.13358349)). We found that in all three cases the inferred radius is very small, around $5$ km, and the NS is extremely compact, even though the mass is only slightly above $1$ $M_{\odot}$. In fact, all these solutions strongly hit our prior upper bound on surface gravity set by the hydrogen atmosphere table limits ($\log_{10}g=15.0$).
While no very obvious patterns were observed in the residuals between the inferred phase-energy resolved signals and the PSR J1231--1411 data (which would indicate a poor fit to the data), we found that the fit to the phase-averaged (or bolometric) NICER pulse profile is rather bad for all the ST-U models (see the left-hand panel of Figure 4 for an example). The bolometric $\chi^{2}$ values for maximum-likelihood solutions are around $80$ for $32$ phase bins. [^24] The degree of freedom (d.o.f.) is not obvious, but it should be a few smaller than $32$ depending on how many of the model parameters influence the bolometric pulse profile (Dittmann et al., 2024). We also note that better bolometric $\chi^{2}$ values are found when looking at the $1000$ highest likelihood samples in each ST-U run because the best bolometric NICER $\chi^{2}$ value does not exactly coincide with the maximum joint (2D) NICER and XMM-Newton likelihood sample. However, these $\chi^{2}$ values are still always above $70$. In addition, if forcing the radius to be above $8$ km (making a full run with the second radius prior case mentioned in Section 3.3.2), both the maximum-likelihood and the corresponding bolometric $\chi^{2}$ values became significantly worse. The former dropped by more than 20 in $\ln$-units and the latter increased to $\sim 160$ (results of this run are available in the Zenodo, Salmi et al. 2024). Most notably, the best-fit model was no longer able to produce the interpulse around phase $0.5$ that is seen in the data (see Figure 1). This was easier for the high compactness star, due to multiple images of the hot spots contributing to the interpulse. The highest Bayesian evidence for a ST-U model (the one including rest of the star emission) was also found to be significantly smaller than for any of the PDT-U models (by $8-33$ in $\ln$-units), so we therefore focus on the latter next.